# USED: Universal Speaker Extraction and Diarization

### v2

@@ III-F Separator
The separator is designed to separate speakers based on the embeddings from the embedding assignment module. As shown in Fig. 1, the embedding sequence $X$ firstly undergoes a series of operations, including a layer normalization followed by $1\times 1$ CNN layer, wherein the input channel size is set to $3C$ and the output channel size is $C$. Then, we introduce $N^{t}$ Temporal Convolutional Network (TCN) blocks [26] to generate output representations based on the embeddings and speech mixture. Representations are then concatenated and down-sampled via a Conv1D layer with input channel size $(k+1)\times C$ and output channel size $C$. Finally, the processed representation is forwarded to another $N^{t}$ TCN blocks to generate representations $\{R_{1},...,R_{{}_{N^{t}}}\}$.
Each TCN block comprises a set of $N^{b}$ TCN layers. For the dilated depth-wise separable convolution (D-Conv) shown in Fig. 2b, it has an exponential growth dilation factor $2^{b}$, where $b\in\{0,...,N^{b}-1\}$. In the initial TCN layer within each block, the input channel size for the $1\times 1$ CNN is $D_{\text{spk}}+C$.
@@ III-H Extraction Decoder
The extraction decoder is designed to estimate masks for each speaker and reconstruct the corresponding signals. As illustrated in Fig. 1, the mask module of extraction decoder comprises a Conv1D and ReLU activation to generate masks $\{M_{1},...,M_{k+1}\}$, where $M_{i}\in T\times C$. We then obtain the modulated responses $i$ by element-wise multiplication of the mask $M_{i}$ and the representations $X$ from speech mixture.
For signal reconstruction, we employ a multi-scale decoder like speech encoder, which consists of three transposed convolutional networks (ConvTran1D), each with a kernel size of $L^{\text{e}}_{1}$, $L^{\text{e}}_{2}$ and $L^{\text{e}}_{3}$, and a common stride, $L^{\text{e}}_{1}/2$. It’s used to reconstruct the time-domain signals, $\tilde{s}=\{\tilde{s}_{1},...,\tilde{s}_{k+1}\}$, where $\tilde{s}_{i}=\{\tilde{s}^{1}_{1},\tilde{s}^{2}_{1},\tilde{s}^{3}_{1}\}$ is from three ConvTran1D layers.
As the USED model generates both the extracted speech and speaker diarization results, we propose an interaction module (IM) to use speaker diarization results to refine the generated waveforms. To avoid the effect of the gradient from the extraction task on the diarization task, we first create a clone of $D^{N^{t}}=\{D^{N^{t}}_{1},...,D^{N^{t}}_{(k+1)}\}$ without gradient. Subsequently, we employ interpolation techniques on the diarization output probabilities to ensure length-based uniformity between the two tasks. Following this alignment, a layer comprising a Conv1D layer and ReLU activation is employed to further process the output probabilities and obtain outputs $r_{i}$, ensuring that the values at each step are not constrained to be less than 1. The Conv1D layer possesses an input channel size of 1, an output channel size of 1, and a kernel size of $L^{\text{im}}$.
Finally, the model generates the extracted speeches, $\{\hat{s}_{1},...,\hat{s}_{k+1}\}$, through element-wise multiplication of $\{r_{1},...,r_{k+1}\}$ and $\tilde{s}$ from extraction module separately as
@@ III-I2 Speaker Extraction
SAD loss is used to calculate the loss between the extracted speeches $\{\hat{s}_{1},..,\hat{s}_{k+1}\}$ generated from the multi-task interaction module and target speeches $\{s_{1},...,s_{k+1}\}$. For clarity, we assume the first $m$ target speeches correspond to the embeddings with active state and are the same as $\{c_{1},...,c_{m}\}$ in Eq. (1). $\{s_{m+1},...,s_{k}\}$ are zero vectors which present silence as the target of embeddings with blank state. $s_{k+1}$ is the target speech for residual prediction. In practice, the order is shuffled as introduced in section III-E. Each pair of extracted speech and target speech are firstly segmented according to four scenarios. Finally, the total speaker extraction loss, i.e. $\mathcal{L}_{\text{ext}}$, is as follows:
@@ III-I3 Speaker Diarization
To optimize the model for the diarization task, we apply the binary cross-entropy loss (BCE), denoted as $\mathcal{L}_{\text{diar}}^{j}$ on the output of $N^{t}$ TCN blocks,
@@ III-I4 Speaker Classification
This speaker encoder is optimized through a cross-entropy loss (CE), denoted as $\mathcal{L}_{\text{spk}}$, for speaker classification during the training process, as
@@ IV-A3 CALLHOME
We use the CALLHOME dataset [49] to evaluate the diarization task with real recordings. The CALLHOME dataset is partitioned into two parts based on the Kaldi recipe  [^3] https://github.com/kaldi-asr/kaldi/tree/master/egs/callhome_diarization/v2. The first part, known as Part 1, is utilized for model adaptation, whereas the second part, referred to as Part 2, is used for evaluation. The CALLHOME dataset consists of telephone-channel recordings with 8k sample rate. Besides CALLHOME Part 1, we further add Libri2Mix 100h, Libri3Mix 100h, Libri2Mix 360h and Libri3Mix 360h, which generates from LibriSpeech [43] train-clean-100 and train-clean-360, as the pre-training dataset. In total, our dataset encompasses approximately 452 hours of data. We then finetune the model only with CALLHOME Part 1 and evaluate on CALLHOME Part 2. During inference, we follow the evaluation pipeline of TS-VAD to evaluate our model so that it does not require pre-enrolled speech. Specifically, we extract the speech references based on the diarization result of spectral clustering and then do inference for USED models. The speaker embedding for spectral clustering is extracted from the ECAPA-TDNN model trained on VoxCeleb2  [^4] https://github.com/TaoRuijie/ECAPA-TDNN.
@@ IV-B Baseline Configuration
As for the TS-VAD baseline, we follow [12] to design the model, with several notable modifications. Firstly, we substitute the bidirectional LSTM in the original TS-VAD model with four transformer layers, with sine-cosine positional encoding as input. The model architecture entails two layers dedicated to processing inputs for each speaker individually, followed by a Conv1D layer for down-sampling. Another two layers are used to combine the information from all speakers. For the transformer, we set the embedding dimension to 384 and employ four attention heads. Second, we use a pre-trained speaker encoder, ECAPA-TDNN, to extract speaker embeddings. This speaker encoder also takes on the role of replacing the CNN encoder to extract time-level representations from the speech mixture. Finally, we use a similar strategy as stated in section III-E, which randomly selects speaker embeddings from the whole dataset. $p_{em}$ is set to be 0.3. The model is optimized using Adam with batch size 64 for 40k steps. The speech encoder is kept fixed for the first 4,000 steps, and the learning rate follows a polynomial decay schedule with an initial value of $2\times 10^{-4}$, which has a warm-up phase for the first 10% of steps. The DER metric is our criterion for selecting the best checkpoint for the TS-VAD model.
@@ IV-C USED Model Configuration
The maximum number of speakers $k$ is 3 for evaluation on the LibriMix and SparseLibriMix datasets and in the assessment on the CALLHOME dataset. For the multi-scale speech encoder, $C$ is 256 and the kernel sizes, $L^{e}_{1}$, $L^{e}_{1}$ and $L^{e}_{1}$, are 20, 80 and 160. The number of ResNet blocks $N^{r}$ in the speaker encoder is set to 4, and the dimension of the speaker embedding $D_{\text{spk}}$ is 256. Regarding the separator module, we set $N^{t}$ to 3 and $N^{b}$ to 8. For the diarization decoder, the CNN layer has a kernel size $L^{\text{diar}}$ of 32 and stride $L^{\text{diar}}/2$ of 16. The mask module of the extraction decoder consists of a CNN with a kernel size of 1 and stride 1, and ReLU activation. For the multi-task interaction module, the kernel size $L^{\text{im}}$ of the CNN layer is 16. The loss weights for outputs from different ConvTran1D layers, $\mu_{1}$, $\mu_{2}$ and $\mu_{3}$ are 0.8, 0.1 and 0.1 by following [7]. We set $0.001$ for $\alpha$ and $\beta$, and $1.0$ for $\gamma$ and $\delta$, which are the loss weights in Eq. (9). $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are set to 1.0. $p_{\text{em}}$ is set to 0.3. $p_{\text{sel}}$ and $p_{\text{res}}$ are set to 0.5 and 0.9, respectively, when residual state is applied. For the USED model without residual state and with $p_{sel}=0$, we call it “USED-F(ix)” since it always gets speech references of present speakers during training.
We optimize the model with Adam on 4 GPUs for 100k steps. The utterances are split into chunks with a size of 4 seconds and a shift of 2 seconds. We set the maximum tokens to 260k, which results in processing around 16 seconds of audio for each batch. The learning rate is set to $31e-$ 3 for all experiments, which is warmed up for the first 10% steps and decayed polynomially for the rest steps. The overall loss value is our primary metric for selecting the best checkpoint for the USED model.
@@
Model Overall Performance SS&SQ QS QQ
@@
IV-D Evaluation Metrics
We use diarization error rate (DER (%)), including speaker confusion (SC (%)), false alarm (FA (%)), and missed detection (MS (%)) to evaluate the speaker diarization performance. Collar tolerance and median filtering are set to 0 seconds and 11 frames for the LibriMix and SparseLibriMix datasets, and 0.25 seconds and 11 frames for the CALLHOME datasets. We report SI-SDR improvement (SI-SDRi (dB)), Power (dB/s), SDR improvement (SDRi (dB)), Short-Time Objective Intelligibility (STOI) [52] and Perceptual Evaluation of Speech Quality (PESQ) [53] for speaker extraction performance.
@@ IV-D Evaluation Metrics
V Experimental Results and Analysis
V-A Results on the LibriMix Dataset
Table I and II show the results on the LibriMix dataset with max mode and min mode, respectively. As indicated before, min mode mainly contains highly overlapped speech, while a larger proportion of max mode is non-overlapped speech compared to min mode. We provide detailed results under different scenarios for the max mode as shown in Table I. Please note that we use the average duration length that anyone is speaking in seconds as the metric of speaker diarization for the QQ scenario.
Our analysis involves comparisons with other popular systems from the literature, including TS-VAD [12], EEND-EDA [11], HuBERT Base [51] and wav2vec 2.0 Base [50] for speaker diarization, SpEx+ [7], and Conv-TasNet [26] for speaker extraction and speech separation, and EEND-SS [23] for joint optimization. For HuBERT Base and wav2vec 2.0 Base, we firstly follow SUPERB diarization task [46] to train two downstream models for two speakers and three speakers, respectively. Subsequently, we combine the output results of these two models for evaluation.
[…]
Model Pre-training Data Speaker Number
2 3 4 5 6 all
AHC clustering [11] - 15.45 18.01 22.68 31.40 34.27 19.43
spectral clustering [39] - 16.05 18.77 22.05 30.46 36.85 19.83
EEND-EDA [11] 15,516h 8.50 13.24 21.46 33.16 40.29 15.29
SC-EEND [54] 15,516h 9.57 14.00 21.14 31.07 37.06 15.75
AED-EEND [55] 15,516h 6.96 12.56 18.26 34.32 44.52 14.22
TS-VAD †[12] 452h 12.95 14.57 20.26 27.36 32.66 15.51
USED-F 452h 9.09 12.76 14.67 26.96 26.71 13.16
USED 452h 10.23 12.78 14.68 32.02 25.10 13.51
@@ V-B Results on SparseLibriMix
In addition to the LibriMix dataset, we are also interested in assessing the performance of our model across a range of overlap ratios. Therefore, we conduct an evaluation on SparseLibriMix, as illustrated in Fig. 4. We select TS-VAD and SpEx+ models as the baseline systems. The model trained on the max mode of Libri2Mix and Libri3Mix is evaluated on SparseLibriMix with 2 and 3 speakers. The proposed USED models outperform each baseline task, confirming our models’ effectiveness for variable values of overlap ratio.
@@ V-C Results on CALLHOME
To compare the performance of diarization on real data, we evaluate our model on the CALLHOME dataset. The results are summarized in Table III, compared with several speaker diarization methods, including two clustering-based methods, AHC clustering [11] and spectral clustering [39], end-to-end speaker diarization models, EEND-EDA [11], SC-EEND [54] and AED-EEND [55], and TS-VAD model [12]. Most of those methods use the simulation data from Swichboard-2, Switchboard Cellular, and NIST Speaker Recognition Evaluation [11], which have a variable number of speakers, ranging from one to five. The total number of hours for the pre-training data is around 15,516 hours for those methods. In contrast, we use much less simulation data, which is from LibriMix and only have 2 or 3 speakers. This results in around 452 hours of pre-training data.
Our USED model achieves around 15% relative improvement on overall DER compared to the TS-VAD baseline and performs better with much less pre-training data compared to other methods. We observe that our USED model is more robust in terms of the speaker numbers of utterances. Results show that although our model only uses simulation data from LibriMix, which has a speech mixture with 2 or 3 speakers, the model still achieves better DER results when the speaker number is larger than 3. We believe this gain is from the design of the embedding assignment module.
V-D Investigation of USED Model for Generating Results of Different Numbers of Speakers During Inference
In this section, we compare the USED-F model and the USED model with baseline systems when generating results for different numbers of speakers during inference, as shown in Table V. When $m$ is 1, USED models perform speaker extraction like SpEx+. Under this condition, USED models and SpEx+ have the same input data for a fair comparison. When $m$ is all, it means we provide all speech references for present speakers, and the embedding is $V_{\text{em}}$ for residual state. Otherwise, we use $V_{\text{res}}$ as the embedding for residual state.
TABLE IV: USED-F model performance under different scenarios when using different loss weights for each scenario of SAD loss with or without Multi-Task interaction module. IM means multi-task interaction module. The results of the first line are from the standard setting.
@@ V-D Investigation of USED Model for Generating Results of Different Numbers of Speakers During Inference
All SS SQ QS & QQ
1.0 1.0 0.001 0.001 12.70 12.00 9.17 -24.0 4.75
1.0 1.0 0 0 12.47 12.02 9.58 55.68 4.80
1.0 1.0 0.001 0.001 12.71 12.00 9.11 -1.93 4.65
1.0 1.0 0.01 0.01 12.50 11.65 8.68 -45.16 4.77
1.0 1.0 1.0 1.0 11.54 10.48 7.42 -79.21 5.57
TABLE V: Model performance on the LbriMix dataset for max mode when setting different values of $m$. RS stands for residual state.
@@
First, USED models consistently perform much better than the SpEx+ baseline when $m$ is 1, except for the USED-F model. This indicates that even if the USED model only gets one speech reference, it can still outperform the SpEx+ model. Second, we observe that the USED model without residual state achieves improvement compared to the USED-F model on both speaker diarization and extraction tasks if $m$ equals 1 or 2. It gets a larger improvement when applying residual state with $p_{\text{res}}$ of 1.0. However, when $m$ is all, the performance of these two models degrades significantly on the speaker diarization task compared to the USED-F model, where the DER increases from 4.75% to 10.38% and 9.46%. This degradation can be avoided by setting $p_{\text{res}}$ to 0.9 during training. In this way, the USED model can achieve a comparable performance compared to the USED-F model if $m$ is all and still significantly outperforms baseline systems when $m$ is 1 or 2.
V-E Ablation Study of USED Models Trained on Single Task
TABLE VI: Ablation study of the USED model with the single task on the LibriMix dataset for max mode.

### v3

@@ III-F Separator
The separator is designed to separate speakers based on the embeddings from the embedding assignment module. As shown in Fig. 1, the embedding sequence $X$ firstly undergoes a series of operations, including a layer normalization followed by $1\times 1$ CNN layer, wherein the input channel size is set to $3C$ and the output channel size is $C$. Then, we introduce $N^{t}$ Temporal Convolutional Network (TCN) blocks [28] to generate output representations based on the embeddings and speech mixture.
We use TCN as the backbone of our model because they have been widely adopted in speech separation and speaker extraction tasks[28, 9]. The USED model is built on the SpEx+ model, which also utilizes TCN. This choice ensures consistency and enhances performance. In addition, TCN is commonly used for time-domain approaches, avoiding the phase estimation issues found in time-frequency models. Finally, compared to methods like DPRNN [29] and BSRNN [30], Conv-TasNet [28] offers lower computational complexity, reduced memory requirements, and faster training and inference.
Representations are then concatenated and down-sampled via a Conv1D layer with input channel size $(k+1)\times C$ and output channel size $C$. Finally, the processed representation is forwarded to another $N^{t}$ TCN blocks to generate representations $\{R_{1},...,R_{{}_{N^{t}}}\}$.
Each TCN block comprises a set of $N^{b}$ TCN layers. The dilated depth-wise separable convolution (D-Conv) shown in Fig. 2b has an exponential growth dilation factor $2^{b}$, where $b\in\{0,...,N^{b}-1\}$. In the initial TCN layer within each block, the input channel size for the $1\times 1$ CNN is $D_{\text{spk}}+C$.
@@ III-H Extraction Decoder
The extraction decoder is designed to estimate masks for each speaker and reconstruct the corresponding signals. The mask module of the extraction decoder comprises a Conv1D layer and ReLU activation to generate masks $\{M_{1},...,M_{k+1}\}$, where $M_{i}\in\mathbb{R}^{T\times C}$. We then obtain the modulated responses $i$ by element-wise multiplication of the mask $M_{i}$ and the representations $X$ from the speech mixture.
For signal reconstruction, we employ a multi-scale decoder like speech encoder, which consists of three transposed convolutional networks (ConvTran1D), each with a kernel size of $L^{\text{e}}_{1}$, $L^{\text{e}}_{2}$ and $L^{\text{e}}_{3}$, and a common stride, $L^{\text{e}}_{1}/2$. It’s used to reconstruct the time-domain signals, $\tilde{s}=\{\tilde{s}_{1},...,\tilde{s}_{k+1}\}$, where $\tilde{s}_{i}=\{\tilde{s}^{1}_{i},\tilde{s}^{2}_{i},\tilde{s}^{3}_{i}\}$ is obtained from three ConvTran1D layers.
As the USED model generates both the extracted speech and speaker diarization results, we propose an interaction module (IM) to use speaker diarization results to refine the generated waveforms. To avoid the gradient from the extraction task having an effect on the diarization task, we first create a clone of $D^{N^{t}}=\{D^{N^{t}}_{1},...,D^{N^{t}}_{(k+1)}\}$ without gradient. Subsequently, we employ interpolation techniques on the diarization output probabilities to ensure length-based uniformity between the two tasks. Following this alignment, a layer comprising a Conv1D layer and ReLU activation is employed to further process the output probabilities and obtain outputs $r_{i}$, ensuring that the values at each step are not constrained to be less than 1. The Conv1D layer possesses an input channel size of 1, an output channel size of 1, and a kernel size of $L^{\text{im}}$.
Finally, the model generates the extracted speeches, $\{\hat{s}_{1},...,\hat{s}_{k+1}\}$, through element-wise multiplication of $\{r_{1},...,r_{k+1}\}$ and the output $\tilde{s}$ from the extraction module separately as
@@ III-I2 Speaker Extraction
SAD loss is used to calculate the loss between the extracted speeches $\{\hat{s}_{1},..,\hat{s}_{k+1}\}$ generated from the multi-task interaction module and target speeches $\{s_{1},...,s_{k+1}\}$. For clarity, we assume the first $m$ target speeches correspond to the embeddings with active state and are the same as $\{c_{1},...,c_{m}\}$ in Eq. (1). $\{s_{m+1},...,s_{k}\}$ are zero vectors which present silence as the target of embeddings with blank state. $s_{k+1}$ is the target speech for residual prediction. In practice, the order is shuffled as introduced in section III-E. Each pair of extracted speech and target speech are firstly segmented according to the four scenarios. Finally, the total speaker extraction loss, i.e. $\mathcal{L}_{\text{ext}}$, is as follows:
@@ III-I3 Speaker Diarization
To optimize the model for the diarization task, we apply the binary cross-entropy loss (BCE), denoted as $\mathcal{L}_{\text{diar}}^{j}$ on the output of the $N^{t}$ TCN blocks,
@@ III-I4 Speaker Classification
The speaker encoder is optimized through a cross-entropy loss (CE), denoted as $\mathcal{L}_{\text{spk}}$, for speaker classification during the training process, as
@@ IV-A3 CALLHOME
We use the CALLHOME dataset [63] to evaluate the diarization task with real recordings. The CALLHOME dataset is partitioned into two parts based on the Kaldi recipe  [^3] https://github.com/kaldi-asr/kaldi/tree/master/egs/callhome_diarization/v2. The first part, known as Part 1, is utilized for model adaptation, whereas the second part, referred to as Part 2, is used for evaluation. The CALLHOME dataset consists of telephone-channel recordings with 8k sample rate. Besides CALLHOME Part 1, we further add Libri2Mix 100h, Libri3Mix 100h, Libri2Mix 360h and Libri3Mix 360h, which are generated from LibriSpeech [58] train-clean-100 and train-clean-360, as the pre-training dataset. In total, our dataset encompasses approximately 452 hours of data. We then finetune the model only with CALLHOME Part 1 and evaluate on CALLHOME Part 2. During inference, we follow the evaluation pipeline of TS-VAD to evaluate our model so that it does not require pre-enrolled speech. Specifically, we extract the speech references based on the diarization result of spectral clustering and then do inference for USED models. The speaker embedding for spectral clustering is extracted from the ECAPA-TDNN model [64] trained on VoxCeleb2 [65]  [^4] https://github.com/TaoRuijie/ECAPA-TDNN.
@@ IV-B Baseline Configuration
As for the TS-VAD baseline, we follow [14] to design the model, with several notable modifications. Firstly, we substitute the bidirectional LSTM in the original TS-VAD model with four transformer layers, with sine-cosine positional encoding as input. The model architecture entails two layers dedicated to processing inputs for each speaker individually, followed by a Conv1D layer for down-sampling. Another two layers are used to combine the information from all speakers. For the transformer, we set the embedding dimension to 384 and employ four attention heads. Second, we use a pre-trained speaker encoder, ECAPA-TDNN [64], to extract speaker embeddings. This speaker encoder also takes on the role of replacing the CNN encoder to extract time-level representations from the speech mixture. Finally, we use a similar strategy as stated in section III-E, which randomly selects speaker embeddings from the whole dataset. $p_{em}$ is set to be 0.3. The model is optimized using Adam with batch size 64 for 40k steps. The speech encoder is kept fixed for the first 4,000 steps, and the learning rate follows a polynomial decay schedule with an initial value of $2\times 10^{-4}$, which has a warm-up phase for the first 10% of steps. The DER metric is our criterion for selecting the best checkpoint for the TS-VAD model.
@@ IV-C USED Model Configuration
The maximum number of speakers $k$ is 3 for evaluation on the LibriMix and SparseLibriMix datasets and in the assessment on the CALLHOME dataset. For the multi-scale speech encoder, $C$ is 256 and the kernel sizes, $L^{e}_{1}$, $L^{e}_{2}$ and $L^{e}_{3}$, are 20, 80 and 160. The number of ResNet blocks $N^{r}$ in the speaker encoder is set to 4, and the dimension of the speaker embedding $D_{\text{spk}}$ is 256. Regarding the separator module, we set $N^{t}$ to 3 and $N^{b}$ to 8. For the diarization decoder, the CNN layer has a kernel size $L^{\text{diar}}$ of 32 and stride $L^{\text{diar}}/2$ of 16. The mask module of the extraction decoder consists of a CNN with a kernel size of 1 and stride 1, and ReLU activation. For the multi-task interaction module, the kernel size $L^{\text{im}}$ of the CNN layer is 16. The loss weights for outputs from different ConvTran1D layers, $\mu_{1}$, $\mu_{2}$ and $\mu_{3}$ are 0.8, 0.1 and 0.1 by following [9]. We set $0.001$ for $\alpha$ and $\beta$, and $1.0$ for $\gamma$ and $\delta$, which are the loss weights in Eq. (9). $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are set to 1.0. $p_{\text{em}}$ is set to 0.3. $p_{\text{sel}}$ and $p_{\text{res}}$ are set to 0.5 and 0.9, respectively, when residual state is applied. The USED model without residual state and with $p_{sel}=0$ is called as “USED-F(ix)” since it always gets speech references of present speakers during training.
We optimize the model with Adam on 4 GPUs for 100k steps, corresponding to approximately 11 epochs. The utterances are split into chunks with a size of 4 seconds and a shift of 2 seconds. We set the maximum tokens to 260k, which results in processing around 16 seconds of audio for each batch. The learning rate is set to $31e-$ 3 for all experiments, which is warmed up for the first 10% steps and decayed polynomially for the remaining steps. The overall loss value is our primary metric for selecting the best checkpoint for the USED model.
IV-D Evaluation Metrics
We use diarization error rate (DER (%)), including speaker confusion (SC (%)), false alarm (FA (%)), and missed detection (MS (%)) to evaluate the speaker diarization performance. Collar tolerance and median filtering are set to 0 seconds and 11 frames for the LibriMix and SparseLibriMix datasets, and 0.25 seconds and 11 frames for the CALLHOME datasets. We report SI-SDR improvement (SI-SDRi (dB)), Power (dB/s), SDR improvement (SDRi (dB)), Short-Time Objective Intelligibility (STOI) [66] and Perceptual Evaluation of Speech Quality (PESQ) [67] for speaker extraction performance. In terms of complexity analysis, we employ Multiply-Accumulate Operations (MACs (G)) to assess the computational cost of the model and perform a calculation of the model’s parameter count.
@@ IV-D Evaluation Metrics
Model Overall Performance SS SQ QS&QQ
@@
V Experimental Results and Analysis
V-A Results on the LibriMix Dataset
Tables I and II show the results on the LibriMix dataset with max mode and min mode, respectively. As indicated before, min mode mainly contains highly overlapped speech, while a larger proportion of max mode is non-overlapped speech compared to min mode. We provide detailed results under different scenarios for the max mode as shown in Table I. Please note that we use the average duration length that anyone is speaking in seconds as the metric of speaker diarization for the QQ scenario.
Our analysis involves comparisons with other popular systems from the literature, including TS-VAD [14], EEND-EDA [13], HuBERT Base [69] and wav2vec 2.0 Base [68] for speaker diarization, SpEx+ [9], and Conv-TasNet [28] for speaker extraction and speech separation, and EEND-SS [25] for joint optimization. For HuBERT Base and wav2vec 2.0 Base, we firstly follow the SUPERB diarization task [61] to train two downstream models for two speakers and three speakers, respectively. Subsequently, we combine the output results of these two models for evaluation.
@@ V-B Results on SparseLibriMix
In addition to the LibriMix dataset, we are also interested in assessing the performance of our model across a range of overlap ratios. Therefore, we conduct an evaluation on SparseLibriMix, as illustrated in Fig. 4. We select TS-VAD and SpEx+ models as the baseline systems. The model trained on the max mode of Libri2Mix and Libri3Mix is evaluated on SparseLibriMix with 2 and 3 speakers. The proposed USED models outperform each task baseline, confirming our models’ effectiveness for variable values of overlap ratio.
@@ V-C Results on CALLHOME
To compare the performance of diarization on real data, we evaluate our model on the CALLHOME dataset. The results are summarized in Table III, compared with several speaker diarization methods, including two clustering-based methods, AHC clustering [13] and spectral clustering [41], PixIT [44], end-to-end speaker diarization models, EEND-EDA [13], SC-EEND [70] and AED-EEND [71], and the TS-VAD model [14]. Most of those methods use the simulation data from Swichboard-2, Switchboard Cellular, and NIST Speaker Recognition Evaluation [13], which have a variable number of speakers, ranging from one to five. The total number of hours for the pre-training data is around 15,516 hours for those methods. In contrast, we use much less simulation data, which is from LibriMix and only have 2 or 3 speakers. This results in around 452 hours of pre-training data.
Our USED model achieves around 15% relative improvement on overall DER compared to the TS-VAD baseline and performs better with much less pre-training data compared to other methods. We observe that our USED model is more robust in terms of the number of speakers. Results show that although our model only uses simulation data from LibriMix, which has a speech mixture with 2 or 3 speakers, the model still achieves better DER results when the number of speakers is larger than 3. We believe this gain is from the design of the embedding assignment module.
@@
Model Pre-training Data Number of Speakers
2 3 4 5 6 all
AHC clustering [13] - 15.45 18.01 22.68 31.40 34.27 19.43
spectral clustering [41] - 16.05 18.77 22.05 30.46 36.85 19.83
PixIT †[44] 79h 17.02 26.03 27.75 36.38 40.77 24.41
EEND-EDA [13] 15,516h 8.50 13.24 21.46 33.16 40.29 15.29
SC-EEND [70] 15,516h 9.57 14.00 21.14 31.07 37.06 15.75
AED-EEND [71] 15,516h 6.96 12.56 18.26 34.32 44.52 14.22
TS-VAD $\ddagger$ [14] 452h 12.95 14.57 20.26 27.36 32.66 15.51
USED-F 452h 9.09 12.76 14.67 26.96 26.71 13.16
USED 452h 10.23 12.78 14.68 32.02 25.10 13.51
TABLE IV: Model performance on the LbriMix dataset for max mode when setting different values of $m$. RS stands for residual state.
@@
V-D Investigation of USED Models for Generating Results of Different Numbers of Speakers During Inference
In this section, we compare the USED-F model and the USED model with baseline systems when generating results for different numbers of speakers during inference, as shown in Table IV. As introduced in Section III-I2, $m$ is the number of embeddings with an active state as the input of the USED model. During inference, $m$ also represents the number of speakers the USED model is required to predict. When $m$ is 1, USED models perform speaker extraction like SpEx+. Under this condition, USED models and SpEx+ have the same input data for a fair comparison. When $m$ is all, it means we provide all speech references for present speakers, and the embedding is $V_{\text{em}}$ for residual state. Otherwise, we use $V_{\text{res}}$ as the embedding for residual state.
First, USED models consistently perform much better than the SpEx+ baseline when $m$ is 1, except for the USED-F model. This indicates that even if the USED model only gets one speech reference, it can still outperform the SpEx+ model. When $m$ is 2 or all, the performance of USED in speaker extraction significantly surpasses that of SpEx+. The experimental results show that the USED model demonstrates superior efficiency and performance in speaker extraction tasks, particularly in multi-speaker scenarios. Unlike SpEx+, which requires multiple inferences for each speaker, USED performs inference just once, reducing redundant computations like those in the speech encoder. It can also simultaneously utilize the embeddings of multiple speakers, resulting in more discriminative results, whereas SpEx+ processes one speaker at a time. As the number of speakers increases, USED’s performance advantage becomes even more significant.
Second, when extracting results for all speakers, the USED model outperforms TS-VAD, as shown in the experimental results. In addition, the TS-VAD model can only extract results for all speakers simultaneously, while the USED model provides flexibility to extract results for a specific speaker or multiple speakers, which is advantageous in scenarios where obtaining embeddings for all speakers at once is not feasible. Furthermore, the USED model does not require a pre-trained speaker encoder, unlike TS-VAD, which typically relies on one.
Third, we observe that the USED model without residual state achieves improvement compared to the USED-F model on both speaker diarization and extraction tasks if $m$ equals 1 or 2. It gets a larger improvement when applying residual state with $p_{\text{res}}$ of 1.0. However, when $m$ is all, the performance of these two models degrades significantly on the speaker diarization task compared to the USED-F model with the DER increasing from 4.75% to 10.38% and 9.46%, respectively. This degradation can be avoided by setting $p_{\text{res}}$ to 0.9 during training. In this way, the USED model can achieve a competitive performance compared to that of the USED-F model if $m$ is all and still significantly outperforms baseline systems when $m$ is 1 or 2.
The performance gap between USED-F and USED models when $m=\text{all}$ is due to the fact that the input to the USED-F model consistently includes speaker embeddings for all speakers during training. As a result, USED-F is designed to predict for all speakers simultaneously, making it specifically optimized for scenarios where $m=\text{all}$. However, its performance is less robust in cases where $m\neq\text{all}$ (e.g., $m=1$ or $m=2$), leading to lower performance compared to the USED model in these situations.
TABLE V: USED-F model performance under different scenarios when using different loss weights for each scenario of SAD loss with or without Multi-Task interaction module. IM means multi-task interaction module. The results of the first line are from the standard setting.
[…]
All SS SQ QS & QQ
0.001 0.001 1.0 1.0 12.70 12.00 9.17 -24.00 4.75
0 0 1.0 1.0 12.47 12.02 9.58 55.68 4.80
0.001 0.001 1.0 1.0 12.71 12.00 9.11 -1.93 4.65
0.01 0.01 1.0 1.0 12.50 11.65 8.68 -45.16 4.77
1.0 1.0 1.0 1.0 11.54 10.48 7.42 -79.21 5.57
TABLE VI: Ablation study of the USED model configured for a single task (Only SD or Only SE) on the LibriMix dataset for max mode.